A candy maker surveyed chocolate bars available in a local supermarket and found the least squares regression model shown below. Use the model to answer parts a through c below.

(Calories) = 28.5 + 10.84Fat(g) + 2.39Sugar(g)

​a) The​ hand-crafted chocolate she makes has 1616g of fat and 2121g of sugar. How many calories does the model predict for a​ serving?
Answer #1

Answer

Regression equation is given as Calories = 28.5 + 10.84Fat(g) + 2.39Sugar(g)

we have to find the number of calories for 16g fat and 21g of sugar

setting the values, fat =16 and sugar =21

this gives us

= 28.5 + (10.84*16) + (2.39*21)

= 28.5 + 173.44 + 50.19

= 252.13  

Number of calories = 252.13 (rounded to 2 decimals) or 252 (rounded to nearest integer number)